RHIC Startup: The work done yesterday to reduce the beta* in the Blue beam went very well. The luminosity has been increased without significantly increasing the backgrounds, at least according to STAR. PHENIX has still to report. Yellow lifetime appears to have deteriorated somewhat. The beam intensities per bunch in the AtR are 1.2E11(Blue) and 1.0E9(Yellow). Previous running had seen 1.4 in both, so there is some room for improvement.
PHENIX recovered the North Muon Arm during an access late Wednesday.
Mike Brennan reports that the stochastic cooling amplifiers are maxed out, showing that there is something wrong and maybe the cooling may not be functioning at all.
Vladimir points out we have not done a 93 bunch store or made use of the fancy gap store yet. He also notes that the Yellow and Blue tunes drift towards each other, This causes emittance blow up. Usually it is the Yellow tune that drifts.
Christoph will incrementally try to improve the efficiency at transition by playing with the chromaticity.
Mike Mapes has flashed the TSPs in the warm section between 8 and 9 oclock in the Blue ring. After the next fill (`16:30) he will flash the TSPs between 7 and 8 oclock, just upstream of the PHENIX Muon ID. If this makes a measurable change in the background levels, then flashing of these TSPs will become part of the routine.
PHENIX and STAR have requested a brief 15 minute access after the 16:30 beam is dumped. This will coincide with the vacuum work.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |